OpenStudy (anonymous):

Please HELP!! -- Explain, in complete sentences, the purpose of converting a radical expression into a rational exponent. Give an example.

OpenStudy (anonymous):

Because then radical expressions with different indices can be multiplied and divided using the laws of exponents.

OpenStudy (anonymous):

\[\sqrt[3]{2}\sqrt[5]{2}=2^{1/3}*2^{1/5}=2^{8/15}\]

OpenStudy (anonymous):

@meanzie

OpenStudy (anonymous):

Thank You So So Much!!!

OpenStudy (anonymous):

anytime :)